A company has just paid its first dividend of $3.30. Next year's dividend is forecast to grow by 7 percent, followed by another 7 per cent growth in year two. From year three onwards dividends are expected to grow by 2.5 percent per annum, indefinitely. Investors require a rate of return of 12 percent p.a. for investments of this type. The current price of the share is (round to nearest cent)
Answer #1

Let current price of stock be P0

Let current period be Period 0, and Dividend in Period 0 = D0

Given D0 = 3.30

D1 = 3.30*1.07 = 3.531

D2 = 3.531*1.07 = 3.778

Hence, P1 = D2/(r-g) = 3.778/(0.12-0.025) = 39.77

P0 = P1/(1+r) = 39.77/1.12 = $35.50